# Visual Whole-Body Control for Legged Loco-Manipulation

### v4

@@ Abstract
We study the problem of mobile manipulation using legged robots equipped with an arm, namely legged loco-manipulation. The robot legs, while usually utilized for mobility, offer an opportunity to amplify the manipulation capabilities by conducting whole-body control. That is, the robot can control the legs and the arm at the same time to extend its workspace. We propose a framework that can conduct the whole-body control autonomously with visual observations. Our approach, namely Visual Whole-Body Control (VBC), is composed of a low-level policy using all degrees of freedom to track the body velocities along with the end-effector position, and a high-level policy proposing the velocities and end-effector position based on visual inputs. We train both levels of policies in simulation and perform Sim2Real transfer for real robot deployment. We perform extensive experiments and show significant improvements over baselines in picking up diverse objects in different configurations (heights, locations, orientations) and environments.
@@
Figure 1: Real-Robot input visualization and grasping trajectory. Our framework enables the robot to grasp different objects in varying heights and surroundings. Our policy is trained in simulation using visual inputs (masks and masked depth images from two views) and can be deployed in real robots without any real-world fine-tuning.
00footnotetext: The first two authors contributed equally.
I Introduction
@@
The study of mobile manipulation has achieved large advancements with the progress of better manipulation controllers and navigation systems. While installing wheels for a manipulator can help solve most household tasks [75, 23], it is very challenging to adopt these robots outdoors with challenging terrains. Imagine going camping; having a robot picking up trash and wood for us would be very helpful. To achieve such flexibility and applications in the wild, we study mobile manipulation with legged robots equipped with an arm, i.e., legged loco-manipulation. Specifically, we are interested in how the robot can conduct tasks based on its visual observation autonomously.
@@
An important aspect of loco-manipulation is the opportunity to conduct whole-body control using all degrees of freedom of the robot. While the legs are typically utilized for mobility, there is a large potential for controlling all the joints of the legs and arms together to amplify manipulation capabilities. We humans do this, too: We use our legs to stand up to reach a book placed on a shelf; We bend our legs to tie our shoes. Fig. 1 shows the tasks our robot can do when controlling the whole-body joints together. For example, our robot can go to the side of the road and pick up the trash on the grass. This will be unachievable without bending the front legs of the robot. The arm alone will not be long enough to reach the ground.
@@
While this is appealing, it is a very challenging control problem to coordinate all the joints simultaneously with large degrees of freedom (19 DoF). The robot will need to exploit the contact with its surroundings and the objects and maintain stability and robustness to external disturbances all at the same time. Besides, acquiring a mobile robot platform to do precise manipulation jobs requires a steady robot body, which is much harder for legged robots compared to wheeled robots. An even larger challenge comes from achieving all these in diverse environments autonomously, given only the observation from an egocentric camera. Recent learning-based approaches have shown promising results on legged robots avoiding obstacles, climbing stairs, and jumping over stages robustly with visual inputs [13, 1, 82, 15]. However, all these efforts still focus only on locomotion without manipulation, which requires more precise control. The extra complexity and more challenging tasks make direct end-to-end learning infeasible.
@@
In this paper, we conduct loco-manipulation by introducing a two-level framework with a high-level policy proposing end-effector pose and robot body velocity commands based on visual observations and a low-level policy tracking these commands. Such a hierarchical design is effective as the low-level controller can be effectively shared when interacting with diverse environments and objects. We named our framework Visual Whole-Body Control (VBC). Specifically, VBC is a hierarchical solution that contains three stages of training: first, we train a universal low-level policy using reinforcement learning (RL) to track any given goals and achieve whole-body behaviors; then, we train a privileged teacher policy by RL to provide a set of appropriate goals and guide the low-level policy to accomplish specific tasks (e.g., pick-up); finally, to deploy the policy into the real robot, we distill the teacher policy into a depth-image-based visuomotor student policy via online imitation learning. All this training is done in simulation, and we perform direct Sim2Real transfer for deployment in the real robot.
@@
Our hardware platform is built on a Unitree B1 quadruped robot equipped with a Unitree Z1 robotic arm. We design a server-client multi-processing system that consists of four different modules working under different frequencies. We conducted the pickup task with 14 different objects, including regular shapes, daily uses, and irregular objects, under three height configurations, i.e., on the ground, on a box, and on a table, achieving a high success rate in all scenarios.
We summarize the contribution of this paper as:
  • We develop an autonomous vision-based loco-manipulation system that utilizes model-free RL and sim-to-real techniques to learn whole-body control policies for legged robots.
  • Our proposed method emerges retrying behavior and generalizes well to complicated terrains, various objects, and heights.
II Related Work
@@
Legged Loco-Manipulation A set of works has exploited legged loco-manipulation to take advantage of legged robots. Bellicoso et al. [5] designs an online motion planning framework, together with a whole-body controller based on a hierarchical optimization algorithm to achieve loco-manipulation, Fu et al. [22] proposed an effective method to train a whole-body controller end-to-end through DRL. Ma et al. [46] combined model-based manipulator control and learning-based locomotion, but only showed the response to external push instead of achieving manipulation tasks. Notably, these works all require human involvement and teleoperation to execute and achieve tasks, rather than operating autonomously. Zhang et al. [81], Yokoyama et al. [79], Arcari et al. [4] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default model-based controller provided through the robot APIs (only control robot locomotion in a static height) and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects. Ferrolho et al. [20] proposed a trajectory optimization framework based on a robustness metric for solving complex loco-manipulation tasks, but the task is assigned and the target positions are calculated via a motion capture system, thus only works in limit. Sleiman et al. [60] realized a bilevel search strategy for motion planning, by incorporating domain-specific rules, combing trajectory optimization, informed graph search, and sampling-based planning, enabling robots to perform complex tasks. They require the user to provide a dense description of the scene to initialize the solver. Zimmermann et al. [83] also utilized trajectory optimization to enable a legged robot to walk and pick up objects. However, they only show limited surroundings and simple objects. Besides these works with arm-based configuration, Wolfslag et al. [67] use the legs for pushing and box lifting tasks by introducing additional support legs named prongs to enhance the robustness and manipulation capabilities of quadrupedal robots. Jeon et al. [32] adopted a similar hierarchical framework to us, but does not involve any visual inputs, and only learned a quadruped for pushing large objects without an additional arm.
@@
The system we design highly differs from the aforementioned works, as our work is a fully autonomous and vision-based mobile manipulation system on quadruped robots, with limited human intervention. The robot is controlled and planned by an end-to-end policy integrated with perception and control, able to adapt to objects with varying heights through its whole-body behavior, and emerges retrying behaviors when the task fails. Due to our design of visual inputs, our system can work both indoors and outdoors with no external limits, as illustrated in Fig. 1, compared with almost all of the above-mentioned works that are only tested in limited lab scenarios. In Fig. 2 we show why VBC is highly advantaged of the non-whole-body methods [81, 79, 4], for the body control behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
(a) Non-whole-body method.
(b) VBC (ours).
Figure 2: Illustration of comparing VBC to non-whole-body method. The robot body of the non-whole-body method does not bend to allow the arm to reach the objects on the ground.
Figure 3: Training a vision-based whole-body policy for loco-manipulation. Our VBC framework trains a low-level control policy and a high-level planning policy using reinforcement learning and imitation learning.
III Visual Whole-Body Control
@@
In this section, we introduce our Visual Whole-Body Control (VBC) framework (as described in Fig. 3). Our VBC framework consists of a low-level goal-reaching policy and a high-level task-planning policy. Our low-level goal-reaching policy tracks a root velocity command and a target end-effector pose. Our vision-based high-level task-planning policy provides velocity and end-effector pose command to the low-level policy, given the segmented depth images and proprioception of the robot as inputs. When combined, our robot could operate fully autonomously using only visual input and its proprioception. We start with an overview of our robot setup, followed by the training pipeline of the low-level policy and the high-level policy, and our robot system design.
III-A Robot Platform Overview
The robot platform used in this work comprises a Unitree B1 quadruped robot with a Unitree Z1 robot arm and a gripper mounted on its back. Our robot is equipped with two cameras, one on the head and another mounted close to the gripper. The Unitree B1 quadruped has 12 actuatable DoFs, the arm has 6 DoFs, and the gripper has another 1 DoF, so there are 19 DoFs in total. An overview of the hardware setup is shown in Fig. 4.
@@ III-A Robot Platform Overview
III-B Low-Level Policy for Whole-Body Goal-Reaching
The low-level control policy that takes over the whole-body locomotion control $\pi^{\text{low}}$ as described in the blue part in Fig. 3, is trained to track any given end-effector poses and body velocities across various terrains. In this work, we turn to the help of RL and utilize ROA (Regularized Online Adaptation) [22, 12, 32] to realize such a universal and robust behavior that can become the foundation of any high-level tasks.
Commands. The command $\mathbf{b}_{t}$ for our low-level policy is defined as:
$\mathbf{b}_{t}=[\mathbf{p}^{\text{cmd}},\mathbf{o}^{\text{cmd}},v_{\text{lin}}^{\text{cmd}},\omega_{\text{yaw}}^{\text{cmd}}]$
[…]
Figure 5: Comparison between trajectories and segmented depth images in simulation and real-world sampled by VBC. Left: The real-world behaviors match great with the simulation, indicating a less sim-to-real gap in our method. Right: Our TrackingSAM-based working pipeline allows us to obtain the segmentation mask precisely during deployment (not the same object observed). The depth images are pre-processed by hole-filling and clipping.
Observations. The observation of our policy is defined as a 90-dimensional vector $\mathbf{o}_{t}$:
$\mathbf{o}_{t}=[\mathbf{s}_{t}^{\text{base}},\mathbf{s}_{t}^{\text{arm}},\mathbf{s}_{t}^{\text{leg}},\mathbf{a}_{t-1},\mathbf{z}_{t},\mathbf{t}_{t},\mathbf{b}_{t}].$
Among them, $\mathbf{s}_{t}^{\text{base}}\in\mathbb{R}^{5}$ is the current quadruped base state including (row, pitch, and quadruped base angular velocities), $\mathbf{s}_{t}^{\text{arm}}\in\mathbb{R}^{12}$ is arm state (position and velocity of each arm joint except the end-effector), $\mathbf{s}_{t}^{\text{leg}}\in\mathbb{R}^{28}$ is leg state (joint position and velocity of each leg joint, and foot contact patterns that are 00 or $1$), $\mathbf{a}_{t-1}\in\mathbb{R}^{12}$ is the last output of policy network, $\mathbf{z}_{t}\in\mathbb{R}^{20}$ is environment extrinsic vector, which is an encoding of environment physics. Besides, following the work of [1], to learn a steady walking behavior (like trotting), we provide some extra timing reference variables $\mathbf{t}_{t}$ to the observation, which is computed from the offset timings of each foot by repeat the gait cycle of the desired symmetric quadrupedal contact pattern.
Actions. Our low-level goal-reaching policy controls the quadruped robot and the mounted arm simultaneously. For the quadruped robot, our low-level policy outputs the target joint angles for all 12 joints of the robot, which is converted to torque command by a PD controller.
Arm control. Previous work [22] has revealed that whole-body RL policy is good at controlling the end-effector position but struggles in controlling the end-effector orientation. Unfortunately, such behaviors are critical for our manipulation tasks. Thereafter, we solve the inverse kinematic (IK) with the pseudoinverse Jacobian method222 More advanced IK solver can be exploited for better performance under certain circumstances, but the pseudoinverse Jacobian method works well enough for the problem studied in this paper. to convert the end-effector pose command into target joint angles for controlling the arm. Formally, IK computes the increment of the current joint position $\Delta\mathbf{\theta}\in\mathbb{R}^{6}$
$\Delta\mathbf{\theta}=J^{T}\left(JJ^{T}\right)^{-1}\mathbf{e},$ (1)
where $J$ is the Jacobian matrix, and $\mathbf{e}\in\mathbb{R}^{6}$ is the difference between the current end-effector pose and the next end-effector pose represented as the position and the Euler angle.
Reward function. Our reward for training the low-level policy for whole-body control mainly consists of four parts: command-following reward, energy reward, alive reward, and phase reward. The command following reward encourages the robot to track and follow the provided commands and explore various whole-body behaviors; the energy reward penalizes energy consumption to enable smooth and reasonable motion; the alive reward encourages the robot not to fail; the phase reward encourages the robot to walk in a steady behavior as mentioned above.
End-effector goal poses sampling. To train a robust low-level controller for goal-reaching, we sample and track various gripper poses. For sampling smooth trajectories of gripper target poses for tracking, we employ a height-invariant coordinate system originating at the robot’s base and randomly sample trajectories on this at every fixed timestep. This ensures the sampled trajectories are always on a smooth sphere and prevents the sampled goals from being affected by the quadruped robot’s height. Details can be found in Appendix II-A.
Domain randomization. When training, we randomize the type of terrains, including flat plane and rough ones. We also randomized the friction between robots and terrains. Besides, we randomized the quadruped’s mass and center of mass. These randomizations, together with our ROA module, contribute to a robust low-level policy.
III-C High-Level Policy for Task Planning
When deploying our policy into the real world, the robot can only perceive the object through depth images to achieve high-frequency control. However, RL with visual observation is quite challenging and requires particular techniques to be efficient [30, 77, 29, 28, 70]. To achieve simple and stable training, we choose to train a privileged state-based policy that can access the shape information of the object (as described in the orange part of Fig. 3) and distill a visuomotor student (as described in the green part of Fig. 3).
C.1. Privileged Teacher Policy
Like the low-level policy, the privileged teacher policy is trained through RL (we also use PPO in this work) with designed task-specific reward functions and domain randomization, with a fixed low-level policy underlying.
Privileged observations. The privileged observations include the encoded shape feature and object poses. We formally define it as a $1094$-dim vector:
$\mathbf{o}_{t}=[\mathbf{z}^{\text{shape}},\mathbf{s}_{t}^{\text{obj}},\mathbf{s}_{t}^{\text{proprio}},\mathbf{v}_{t}^{\text{base vel}},\mathbf{a}_{t-1}],$
where $\mathbf{z}^{\text{shape}}\in\mathbb{R}^{1024}$ is the latent shape feature vector encoded from the object point clouds using a pre-trained PointNet++ [55]. The pre-trained PointNet++ is fixed, so $\mathbf{z}^{\text{shape}}$ remains invariant during training. $\mathbf{s}_{t}^{\text{obj}}\in\mathbb{R}^{6}$ is the object pose in local observation w.r.t the robot arm base. It contains the local pose of the object. The use of local coordination inherently incorporates the information regarding the distance to the object. $\mathbf{s}_{t}^{\text{proprio}}\in\mathbb{R}^{53}$ consists of joint positions including gripper joint position $\mathbf{q}_{t}\in\mathbb{R}^{19}$, joint velocity not including gripper joint velocity $\dot{\mathbf{q}}_{t}\in\mathbb{R}^{18}$, and end-effector position and orientation $\mathbf{s}^{\text{ee}}_{t}\in\mathbb{R}^{6}$. $\mathbf{v}_{t}^{\text{base vel}}\in\mathbb{R}^{3}$ is the base velocity of robot and $\mathbf{a}_{t-1}\in\mathbb{R}^{9}$ is the last high-level action.
High-level Actions. The high-level policy $\pi^{\text{high}}$ outputs actions that determine the velocities and the target end-effector position for the low-level policy. The high-level action is formally defined as:
$\mathbf{a}_{t}=[\mathbf{p}_{t}^{\text{cmd}},\mathbf{v}_{t}^{\text{cmd}},p_{t}^{\text{gripper}}]\in\mathbb{R}^{9},$
where $\mathbf{p}_{t}^{\text{cmd}}\in\mathbb{R}^{6}$ is the increment of gripper pose, $\mathbf{v}_{t}^{\text{cmd}}\in\mathbb{R}^{2}$ is the command of quadruped linear velocity and yaw velocity and $p_{t}^{\text{gripper}}\in\{0,1\}$ indicates the gripper status (close or open).
Reward functions. The reward function for training our state-based policy comprises stage rewards and assistant rewards, as defined in Tab. III. As for the task-specific rewards, we design three stages. The first stage is approaching, and the corresponding reward $r_{\text{approach}}$ encourages the gripper and quadruped to get close to the target object:
$r_{\text{approach}}=\min(d_{\text{closest}}-d,0)~{},$ (2)
where $d_{\text{closest}}$ is the current minimum distance between the gripper and the target object. The second stage is task progress, and the reward $r_{\text{progress}}$ leads the robot to the desired final state, e.g., lifting the object in the pickup task:
$r_{\text{progress}}^{\text{pickup}}=\min(d-d_{\text{highest}},0)~{},$ (3)
where $d_{\text{highest}}$ is the current largest height of the object. The third stage is denoted task completion, and reward $r_{\text{completion}}$ bonus when the agent completes the task.
$r_{\text{completion}}=\mathbf{1}(\text{Task is completed})~{},$ (4)
For example, in the pickup task, the task completion condition is that the object height exceeds the required threshold. Note that at one time, the robot can only be at one stage and receive one stage reward of the above three, ensuring that different task stages are not intervened in. The assistant rewards are designed to a) smooth the robot’s behavior to be easier to transfer to the real world, like penalty huge action commands and joint acceleration; b) prevent deviation and sampling useless data, such as acquiring the robot’s body and gripper to point to the object, which guarantees that the camera does not lose tracking objects. Beyond these major rewards, we also design a set of assistant reward functions that help accelerate the learning procedure, see details in Appendix II-B.
@@ C.1. Privileged Teacher Policy
Domain randomization. When training our high-level policy, we randomized the friction between the robot and the terrain, the mass of the robot, and the center of mass of the robot. Although our low-level control policy can handle these gaps in the real world, its performances differ under different environmental conditions. These randomizations help the high-level policy adjust better to different low-level performances. Moreover, due to the uncertainty of the high-low signal transmission in the real world, we also randomize the high-level frequency by involving different low-level policy calls during one high-level call. We also randomize the Kp and Kd parameters for the arm controller as we use the built-in position control mode for the arm in the simulation. Regarding the pick-up task, we randomize table heights, the initial position and pose of the robot, along with the position and pose of objects.

### v5

@@ Abstract
We study mobile manipulation using legged robots equipped with an arm, namely legged loco-manipulation. The robot legs, while usually utilized for mobility, offer an opportunity to amplify the manipulation capabilities by coordinating arms for whole-body control. We propose a framework that can conduct the whole-body control autonomously with visual observations. Our approach, namely Visual Whole-Body Control (VBC), is composed of a low-level control policy using all degrees of freedom to track the body velocities along with the end-effector position, and a high-level task-planning policy proposing the velocities and end-effector position based on visual inputs. We train all policies in simulation and perform Sim2Real transfer for real robot deployment. Extensive experiments show clear advantages over baselines in picking up diverse objects in various configurations (heights, locations, orientations) and environments.
@@
Figure 1: Real-Robot input visualization and grasping trajectory. Our framework enables the robot to grasp different objects in varying heights and surroundings. Our policy is trained in simulation using visual inputs (masks and masked depth images from two views) and can be deployed in real robots without any real-world data collection or fine-tuning.
Keywords: Robot Learning, Whole-Body Control, Legged Loco-Manipulation
1 Introduction
@@
The study of mobile manipulation has achieved large advancements with the progress of better manipulation controllers and navigation systems. While installing wheels for a manipulator can help solve most household tasks [1, 2], it is very challenging to adopt these robots outdoors with challenging terrains. Imagine going camping, having a robot picking up trash and wood for us would be very helpful. To achieve such flexibility and applications in the wild, we study mobile manipulation with legged robots equipped with an arm, i.e., legged loco-manipulation. Specifically, we are interested in how the robot can conduct tasks using visual observation autonomously.
@@
An important aspect of loco-manipulation is the opportunity to use all degrees of freedom (e.g., 19 DoF in this paper) of the robot for whole-body control. While the legs are typically utilized for mobility, there is a large potential for coordinating legs and arms together to amplify manipulation capabilities and workspace. We humans do this, too: we stand up to reach books placed on a shelf; we bend to tie our shoes. Fig. 1 shows tasks our robot can do by such a joint control. For example, our robot can go to the side of the road and pick up the trash on the grass. This will be unachievable without coordinating the legs of the robot, as the arm alone can hardly reach the ground.
@@
While appealing, it is a very challenging control problem. The robot will need to exploit the contact with its surroundings and the objects and maintain stability and robustness to external disturbances all at the same time. Besides, acquiring a mobile platform to do precise manipulation jobs requires a steady robot body, which is much harder for legged robots compared to wheeled ones. An even larger challenge comes from achieving all these in diverse environments autonomously, given only the observation from an egocentric camera. Recent visual-learning-based approaches have shown promising results of locomotion on legged robots, like avoiding obstacles, climbing stairs, and jumping over stages [3, 4, 5, 6]. However, loco-manipulation requires more precise control, and the extra complexity can make direct end-to-end learning infeasible.
@@
In this paper, we introduce a two-level framework for loco-manipulation, where a high-level policy proposes end-effector pose and robot body velocity commands based on visual observations and a low-level policy tracks these commands. Such a low-level controller is general for different tasks and the high-level policy is task-relevant. We named our framework Visual Whole-Body Control (VBC). Specifically, VBC contains three stages of training: first, we train a universal low-level policy using reinforcement learning (RL) to track any given goals and achieve whole-body behaviors; then, we train a privileged teacher policy by RL to provide immediate goals and guide the low-level policy to accomplish a task (e.g., pick-up); finally, we distill the teacher policy into a depth-image-based visuomotor student policy via online imitation learning. All training is done in simulation, and we perform Sim2Real transfer to deploy policies on real robots with zero human-collected data.
@@
Our hardware platform is built on a Unitree B1 quadruped robot equipped with a Unitree Z1 robotic arm. We validate the effectiveness of each designed module of VBC in simulation on various objects and heights. As for the real-world experiments, we conducted the pickup task with 14 different objects, including regular shapes, daily uses, and irregular objects, under three height configurations, i.e., on the ground, on a box, and on a table, achieving a high success rate with emergent retrying behaviors and generalization ability in all scenarios.
2 Related Work
@@
A set of works has exploited legged loco-manipulation to take advantage of legged robots. Bellicoso et al. [7] designs an online motion planning framework, together with a whole-body controller based on a hierarchical optimization algorithm to achieve loco-manipulation. Fu et al. [8] proposed an effective method to train a whole-body controller end-to-end through RL. Ma et al. [9] combined model-based manipulator control and learning-based locomotion, but only showed the response to external push instead of achieving manipulation tasks. Notably, these works all require human involvement and teleoperation to execute and achieve tasks, rather than operating autonomously. Zhang et al. [10], Yokoyama et al. [11], Arcari et al. [12] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default model-based controller provided through the robot APIs (only control robot locomotion in a static height) and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects. Ferrolho et al. [13] proposed a trajectory optimization framework based on a robustness metric for solving complex loco-manipulation tasks, but the task is assigned and the target positions are calculated via a motion capture system, thus only works in limit. Sleiman et al. [14] realized a bilevel search strategy for motion planning, by incorporating domain-specific rules, combing trajectory optimization, informed graph search, and sampling-based planning, enabling robots to perform complex tasks. They require the user to provide a dense description of the scene to initialize the solver. Zimmermann et al. [15] also utilized trajectory optimization to enable a legged robot to walk and pick up objects. However, they only show limited surroundings and simple objects. Besides these works with arm-based configuration, Wolfslag et al. [16] use the legs for pushing and box lifting tasks by introducing additional support legs named prongs to enhance the robustness and manipulation capabilities of quadrupedal robots. Jeon et al. [17] adopted a similar hierarchical framework to us, but does not involve any visual inputs, and only learned a quadruped for pushing large objects without an additional arm.
(a) Static-height.
(b) VBC.
Figure 2: Illustration of comparing VBC to static-height methods. Static-height methods cannot reach objects on the ground without arm-leg coordination.
@@
The system we design highly differs from the aforementioned works, as our work is a autonomous and vision-based loco-manipulation system on quadruped robots, with very limited human annotation before start. The robot is controlled and planned by end-to-end policies integrated with perception and control, able to adapt to objects with varying heights through its whole-body behavior, and emerges retrying behaviors when fails. Due to our design of visual inputs, our system can work both indoors and outdoors without external limits, as illustrated in Fig. 1, compared with almost all of the above-mentioned works that are only tested in limited scenarios. Fig. 2 shows why VBC is highly advantaged of static-height methods [10, 11, 12], for the body behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
3 Visual Whole-Body Control
@@
The proposed VBC framework (Fig. 3) consists of a low-level goal-reaching policy and a high-level task-planning policy. The low-level policy tracks a root velocity command and a target end-effector pose. The high-level policy provides these commands to the low-level policy, based on segmented depth images and proprioception states. When combined, our robot can achieve loco-manipulation tasks autonomously. This section draws details of the proposed framework.
3.1 Low-Level Policy for Whole-Body Goal-Reaching
The low-level control policy $\pi_{\text{low}}$ that takes over the whole-body control as described in the blue part in Fig. 3, is trained to track any given end-effector pose and body velocity across various terrains. We train such a policy with RL to realize a universal and robust behavior that can become the foundation of any high-level tasks.
[…]
Observations. The observation of our policy is formalized as a 90-dimensional vector $\mathbf{o}_{t}=[\mathbf{s}_{t}^{\text{base}},\mathbf{s}_{t}^{\text{arm}},\mathbf{s}_{t}^{\text{leg}},\mathbf{a}_{t-1},\mathbf{z}_{t},\mathbf{t}_{t},\mathbf{b}_{t}].$ Among them, $\mathbf{s}_{t}^{\text{base}}\in\mathbb{R}^{5}$ is the current quadruped base state including row, pitch, and yaw velocities, $\mathbf{s}_{t}^{\text{arm}}\in\mathbb{R}^{12}$ is arm state (position and velocity of each arm joint except the end-effector), $\mathbf{s}_{t}^{\text{leg}}\in\mathbb{R}^{28}$ is leg state (joint position and velocity of each leg joint, and foot contact patterns), $\mathbf{a}_{t-1}\in\mathbb{R}^{12}$ is the last output of policy network, $\mathbf{z}_{t}\in\mathbb{R}^{20}$ is environment extrinsic vector, which encodes environment physics [8]. Following the work of Agarwal et al. [4], we aim to learn a steady walking behavior, such as trotting. To achieve this, we incorporate additional timing reference variables, denoted as $\mathbf{t}_{t}$, into the observation. These variables are computed from the offset timings of each foot by repeating the gait cycle of the desired symmetric quadrupedal contact pattern.
Actions. Our low-level goal-reaching policy mainly controls the quadruped robot. For the quadruped robot, our low-level policy outputs the target joint angles for all 12 leg joints, which is converted to torque command by a PD controller.
Arm control. Previous work [8] has revealed that whole-body RL policy struggles in controlling accurate end-effector position and orientation by controlling arm joint positions, which is critical for our manipulation tasks. Hence, we predict the end-effector pose command and convert it into target joint angles by solving the inverse kinematic (IK) with the pseudoinverse Jacobian method222 Other advanced IK solvers can be exploited for better performance under certain circumstances, but the pseudoinverse Jacobian method works well enough for the problem studied in this paper.. Formally, IK computes the increment of the current joint position $\Delta\mathbf{\theta}\in\mathbb{R}^{6}:$ $\Delta\mathbf{\theta}=J^{T}\left(JJ^{T}\right)^{-1}\mathbf{e},$ where $J$ is the Jacobian matrix, and $\mathbf{e}\in\mathbb{R}^{6}$ is the difference between the current end-effector pose and the next end-effector pose represented as the position and the Euler angle.
End-effector goal poses sampling. To train a robust low-level controller, we sample and track various gripper poses. To make sure we can sample smooth trajectories of gripper target poses for tracking, we employ a height-invariant coordinate system originating at the robot’s base and randomly sample goals in this frame at every fixed timestep. This helps to learn a whole-body behavior and enables an expanded workspace. Details can be referred to Fig. 9 in Appendix C.1.
Due to the page limit, we leave more details of low-level policy training (rewards, policy architecture, goal sampling, randomization, etc.) in Appendix C.1.
Figure 3: The proposed VBC framework for loco-manipulation. VBC trains a low-level control policy and a high-level planning policy using reinforcement learning and imitation learning.
3.2 High-Level Policy for Task Planning
When deploying our policy into the real world, the robot only perceives the object through depth images to achieve high-frequency control. However, RL with visual observation is quite challenging and requires particular techniques to be efficient [18, 19, 20, 21, 22]. To achieve simple and stable training, we train a privileged state-based policy that can access the shape and pose information of the object (the orange part of Fig. 3) and distill a visuomotor student (the green part of Fig. 3).
3.2.1 Privileged Teacher Policy
Like the low-level policy, the privileged teacher policy is trained through RL using task-specific reward functions and domain randomization, with a fixed pre-trained low-level policy underlying.
Privileged observations. The privileged observations include the latent shape features and object poses. We formally define it as a $1094$-dim vector: $\mathbf{o}_{t}=[\mathbf{z}^{\text{shape}},\mathbf{s}_{t}^{\text{obj}},\mathbf{s}_{t}^{\text{proprio}},\mathbf{v}_{t}^{\text{base vel}},\mathbf{a}_{t-1}],$ where $\mathbf{z}^{\text{shape}}\in\mathbb{R}^{1024}$ is the latent shape feature vector encoded from the object point clouds using a pre-trained PointNet++ [23], which is fixed and $\mathbf{z}^{\text{shape}}$ remains invariant during training. $\mathbf{s}_{t}^{\text{obj}}\in\mathbb{R}^{6}$ is the object pose in local observation w.r.t the robot arm base. It contains the local pose of the object. The use of local coordination inherently incorporates the information regarding the distance to the object. $\mathbf{s}_{t}^{\text{proprio}}\in\mathbb{R}^{53}$ consists of joint positions including gripper joint position $\mathbf{q}_{t}\in\mathbb{R}^{19}$, joint velocity not including gripper joint velocity $\dot{\mathbf{q}}_{t}\in\mathbb{R}^{18}$, and end-effector position and orientation $\mathbf{s}^{\text{ee}}_{t}\in\mathbb{R}^{6}$. $\mathbf{v}_{t}^{\text{base vel}}\in\mathbb{R}^{3}$ is the base velocity of robot and $\mathbf{a}_{t-1}\in\mathbb{R}^{9}$ is the last high-level action.
High-level Actions. The high-level policy $\pi^{\text{high}}$ outputs actions that determine the velocities and the target end-effector position command for the low-level policy: $\mathbf{a}_{t}=[\mathbf{p}_{t}^{\text{cmd}},\mathbf{v}_{t}^{\text{cmd}},p_{t}^{\text{gripper}}]\in\mathbb{R}^{9},$ where $\mathbf{p}_{t}^{\text{cmd}}\in\mathbb{R}^{6}$ is the increment of gripper pose, $\mathbf{v}_{t}^{\text{cmd}}\in\mathbb{R}^{2}$ is the command of quadruped linear velocity and yaw velocity and $p_{t}^{\text{gripper}}\in\{0,1\}$ indicates the gripper status (close or open).
@@ 3.2.1 Privileged Teacher Policy
Domain randomization. When training our high-level policy, we randomized the friction between the robot and the terrain, the mass, and the center of mass of the robot. Although the low-level policy already handles some of these gaps, these randomizations help the high-level policy adjust well to different low-level behaviors. Moreover, due to the uncertainty of the high-low signal transmission in the real world, we also randomize the high-level frequency by involving random low-level policy calls during one high-level call. We also randomize the Kp and Kd parameters for the arm controller. Regarding the pick-up task, we randomize table heights, the initial position and pose of the robot, along with the position and pose of objects.
3.2.2 Visuomotor Student Policy